Write a quadratic equation having the given numbers as solutions.
2/3, 9/5

Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

Write a quadratic equation having the given numbers as solutions.
2/3, 9/5
x = 2/3 or x = 9/5
3x-2 = 0 or 5x-9 = 0
---------------------------
Multiply the two equations to get:
(3x-2)(5x-9) = 0
------------------------
Let the expression be an equation that equals zero when x = 2/3 or x = 9/5.
Let y = (3x-2)(5x-9)
Then y = 15x^2-37x + 18
===============================
Cheers,
Stan H.
===============================

Answer by checkley77(12844)

X^2+37X/15+1.2 ANS.